OpenStudy (anonymous):

how do you use the precise definition of limits to prove that the lim x--> 4+ (4/(sqrt(x-4)))=infinity? answer asap please!! :) @Calculus1

OpenStudy (anonymous):

\[\lim_{x \rightarrow 4+}4/\sqrt{x-4}=\infty\]

OpenStudy (anonymous):

also what is \[\lim_{x \rightarrow \infty}x ^{2}(\sqrt{x ^{4}+5}-x ^{2})\]

OpenStudy (anonymous):

when you put 4 instead x it will be \[4/\sqrt{4-4}\] \[4/\sqrt{0} =4/0 = \infty\] \[number /0 =\infty\]